OpenStudy (anonymous):

Find the slope of the line with the given equation. 2x + 4y = 12

OpenStudy (katrinakaif):

Can you turn that equation into a y = mx + b format?

OpenStudy (katrinakaif):

I guess not :P 2x + 4y = 12 4y = -2x + 12 y = -2/4x + 12/4 y = -1/2x + 3

OpenStudy (anonymous):

I was just wondering I f the answer I got was right I got -1/2

OpenStudy (katrinakaif):

Right! ;)

OpenStudy (anonymous):

thanks

OpenStudy (katrinakaif):

Anytime ;)